When Matthew is 12 years old his mother is 3 times as old. How many years must pass before his mother is twice as old? How many years ago was she 7 times as old?
Answer by ankor@dixie-net.com(22740)

When Matthew is 12 years old his mother is 3 times as old.
Matt = 12, Mother = 36
:
How many years must pass before his mother is twice as old?
let y = no. of years for this to be true
y + 36 = 2(y+12)
y + 36 = 2y + 24
36 - 24 = 2y - y
12 = y
This will be true in 12 yrs (Mom will be 48, Matt will be 24)
:
How many years ago was she 7 times as old?
let z = no. of yrs for this to be true
36-z = 7(12-z)
36 - z = 84 - 7z
-z + 7z = 84 - 36
6z = 48
z = 8 yrs for this to be true. (Mom will be 28, Matt will be 4)